# Energy landscape analysis based on the Ising model: Tutorial review

### v1

@@ 2.7 Features of energy landscapes
The disconnectivity graph and basin graph are visualizations of the energy landscape. They are useful for us to intuitively understand the dynamics generating the multivariate time series data and differences between different data or between different groups of data (e.g., between as a group of healthy individuals and a group of those with disease) in terms of the energy landscape. However, we need to quantify such differences to draw conclusions. We refer to such a quantity as a feature of the energy landscape. A primary usage of a feature is to look for association between the feature and a property of interest in the research domain. For example, a schizophrenia researcher may be interested in seeking correlation between the severity of schizophrenia and a feature of the energy landscape. As another example, one may want to measure the effect of treatment in terms of the change in the value of a feature, especially if the feature is known to correlate with the severity of the disease.
@@
    A variant of this feature is an average over different local minima. In [32], the authors used the average of branch length over all the major local minima in the pruned disconnectivity graph. By the pruned disconnectivity graph, we mean one created by sequentially removing insignificant local minima, to keep only the major local minima (see section 3.2.2 for the procedure). The branch length for local minimum $\bm{\alpha}_{i}$ is $\min_{j\text{ such that }j\neq i}\{\overline{E}_{i,j}-E(\bm{\alpha}_{i})\}$, where $\overline{E}_{i,j}$ now refers to the smallest energy at which the $i$th and $j$th branches of the pruned disconnectivity graph merge. For example, in Fig. 3(e), suppose that the pruning process discards $\bm{\alpha}_{2}$, $\bm{\alpha}_{3}$, and $\bm{\alpha}_{6}$, leaving $\bm{\alpha}_{1}$, $\bm{\alpha}_{4}$, and $\bm{\alpha}_{5}$ as major local minima. This pruning is equivalent to setting the threshold energy to the one shown by the dotted horizontal line in the figure to define three major states, each of which is represented by a major local minimum. Then, the branch length for $\bm{\alpha}_{1}$, $\bm{\alpha}_{4}$, and $\bm{\alpha}_{5}$ are equal to $\overline{E}_{1,4}-E(\bm{\alpha}_{1})$, $\overline{E}_{1,4}-E(\bm{\alpha}_{4})$, and $\overline{E}_{1,5}-E(\bm{\alpha}_{5})$, respectively. The last one is equal to $\overline{E}_{4,5}-E(\bm{\alpha}_{5})$ because $\bm{\alpha}_{1}$ and $\bm{\alpha}_{4}$ merge earlier than $\bm{\alpha}_{5}$ does as we climb up the disconnectivity graph from a low energy value. The average branch length used in [32] is the average of these three branch lengths and represents how distinct the different major local minima are from each other.
@@ 3.1 Phase diagram methods
Figure 10: Schematic of the phase diagram of the SK model. A more common parameter in physics literature is the temperature. However, in this figure, we chose the average and standard deviation of $J_{ij}$ because their interpretation is easier in our context.
The statistical mechanical theory exploited here is that of prototypical spin systems, such as the Sherrington-Kirkpatrick (SK) model [67]. The SK model is a type of Ising model and defined by drawing each $J_{ij}$ $(=J_{ji};1\leq i\neq j\leq N)$ independently from the Gaussian distribution with a given mean and standard deviation. The SK model, as well as other spin system models, shows different phases depending on the parameter values (see Fig. 10 for a schematic). In the paramagnetic phase, the spins (i.e., $\sigma_{i}$s) tend to be random and orderless in the sense that the mean activity
@@
Criticality, or phase transition, i.e., the boundary between (typically) the paramagnetic phase and the spin-glass phase in a parameter space (see Fig. 10), is of special interest due to rich dynamical behavior of the spin system near there. On these grounds, Ezaki et al. developed a data-driven method to locate a given multivariate time series data as a point in the phase diagram to tell which phase the given data resides in, and quantify how close the data is from the boundary with a different phase [19].
@@
where $\hat{\mu}$, for example, is the value of $\mu$ for the originally estimated PMEM. With Eq. (17), if we arbitrarily choose the values of $\mu$ and $s$, then we obtain $\bm{J}$, which specifies one PMEM. For each $\bm{J}$ specified by a $(\mu,s)$ pair, we perform Monte Carlo simulations to produce $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ and calculate observables such as $m$ and $q$. In this manner, we can generate a phase diagram for each observable in terms of $\mu$ and $s$, as in Fig. 10.
@@ 3.2.2 Randomization
In Fig. 3, local minimum $\bm{\alpha}_{2}$ is shallow in the sense that the energy barrier between $\bm{\alpha}_{2}$ and $\bm{\alpha}_{3}$ is small. Therefore, the dwell time in $\bm{\alpha}_{2}$ would be small. Therefore, although $\bm{\alpha}_{2}$ is technically a local minimum, it may be better to ignore it in practice. Only the local minima that have sufficiently high energy barriers with others are trustable. We already discussed this topic; by choosing a threshold energy level, shown by the dotted line in Fig. 3(e), one can classify the local minima into groups. However, how should we choose the threshold? How can one tell whether or not a local minimum is too shallow to be seriously considered? This decision has been rather arbitrary in most studies using the ELA. The root of the problem is basically the same as an often asked problem for clustering algorithms for data: how to determine the number of clusters?
@@
We compute the maximal branch length for a sufficiently large number, $c$, of completely random multivariate time series data. We set $c=10^{2}$ in [32] and use $c=10^{3}$ in the following running example in this section. We denote the mean and standard deviation of the maximal branch length over the $c$ random disconnectivity graphs by $\langle L\rangle$ and $\text{std}(L)$, respectively. Then, we remove the local minimum with the shortest branch length in the original disconnectivity graph as being insignificant if its branch is shorter than $\langle L\rangle+2\times\text{std}(L)$. For example, Fig. 7(b) shows the disconnectivity graph for a random multivariate binary time series having the same $N$ ($=7$) and $t_{\max}$ ($=3,272$) as those for the original data whose disconnectivity graph is shown in Fig. 7(a). Due to the randomization, the disconnectivity graph in Fig. 7(b) has shorter branch lengths than the disconnectivity graph in Fig. 7(a) in general. The maximal branch length in Fig. 7(b) is that of $\bm{\alpha}^{\prime}_{3}$ and equal to $L=0.125$. Based on $c=10^{3}$ random disconnectivity graphs, we obtained $\langle L\rangle=0.0902$ and $\text{std}(L)=0.0471$. Therefore, $\langle L\rangle+2\times\text{std}(L)=0.184$. The shortest branch of the empirical disconnectivity graph shown in Fig. 7(a), which is for $\bm{\alpha}_{3}$, is shorter than $0.184$. Therefore, we regard $\bm{\alpha}_{3}$ as an insignificant local minimum and prune it away.
@@
If we have removed a local minimum according to this procedure, then we recompute the branch length of the local minima affected by the removal. In Fig. 7(a), removal of the branch of $\bm{\alpha}_{3}$ does not affect the branch length of any other local minima, especially because $\bm{\alpha}_{4}$ and $\bm{\alpha}_{5}$ merge before they merge $\bm{\alpha}_{3}$. Therefore, there is no need to recompute any branch length. After the possible recomputation, we repeat the above procedure. In other words, if the shortest branch is shorter than $\langle L\rangle+2\times\text{std}(L)$, we remove the branch and recompute the affected branch lengths. For example, in Fig. 7(a), we next remove the branch of $\bm{\alpha}_{2}$ because it is shorter than the threshold (i.e., $0.184$). This removal changes the branch length for $\bm{\alpha}_{1}$ from 0.814 to 1.082. Next, we remove the branch of $\bm{\alpha}_{4}$, which increases the branch length for $\bm{\alpha}_{5}$ from $0.761$ to 0.902.
@@
We repeat the procedure until the branches of all the surviving local minima are longer than the threshold in the pruned disconnectivity graph. We refer to the surviving local minima as major local minima. For example, for Fig. 7(a), the final, pruned discoonnectivity graph, shown in Fig. 7(c), only consists of $\bm{\alpha}_{1}$ and $\bm{\alpha}_{5}$, which are the two major local minima. This pruning algorithm also tells us how to associate each removed local minimum and its basin to a major state. In Fig. 7(a), $\bm{\alpha}_{1}$, $\bm{\alpha}_{2}$, and their basins constitute a major state represented by $\bm{\alpha}_{1}$. Similarly, $\bm{\alpha}_{3}$, $\bm{\alpha}_{4}$, $\bm{\alpha}_{5}$, and their basins constitute a major state represented by $\bm{\alpha}_{5}$.
@@ 3.2.3 Random-walk simulations
where $T_{\bm{\sigma}\to\bm{\sigma}^{\prime}}$ is the transition probability of the random walk from $\bm{\sigma}$ to $\bm{\sigma}^{\prime}$. This detailed balance condition is a quantitative formulation of the already made claim that the movement of the state uphill on the energy landscape is not much likely to occur, especially when the energy barrier to be overcome to go to a different activity pattern is high. In this way, moves from a high energy activity pattern to a low energy one are more likely than vice versa. However, the random walk is a stochastic process, so moves from a lower energy activity pattern to a higher one do occur with some probability. Note that Eq. (18) is consistent with the stationary density given by Eq. (6).
@@ 3.3 Challenges
First, ELA is data-hungry. If we aim at maintaining a high accuracy of fit of the PMEM, we roughly need to increase the data length, $t_{\max}$, twice as we add one variable (i.e., as we increase $N$ by $1$). Therefore, $t_{\max}$ needed increases exponentially in terms of $N$, i.e., in proportion to $2^{N}$ [54]. Therefore, if we seriously take the accuracy, one cannot go beyond approximately $N=10$ or $N=15$, depending on $t_{\max}$ of the data, while ignoring the fitting accuracy to conduct a large-scale ELA (e.g., see section 3.1) is a methodological choice.
@@ 3.4 Variational Bayes approximation method
is an $M\times M$ matrix, and $\rm{diag}(\cdot)$ represents the diagonal matrix whose entries are given by the arguments. In Eq. (23),
@@
$\bm{\bar{\sigma}}\equiv(\sigma_{1},\sigma_{2},\ldots,\sigma_{N},\sigma_{1}\sigma_{2},\sigma_{1}\sigma_{3},\ldots,\sigma_{N-1}\sigma_{N})^{\top}$ (27)
when the PMEM parameters $(h_{1},h_{2},\ldots,h_{N},J_{12},J_{13},\ldots,J_{N-1,N})$ are given by $\bm{\eta}$; matrix $\rm C_{\bm{\eta}}$ is the $M\times M$ covariance matrix of $\bm{\bar{\sigma}}$ when the PMEM parameters are $\bm{\eta}$. In Eq. (24), $\bm{c}_{\bm{\eta}}$ is the vector composed of the diagonal entries of $\rm C_{\bm{\eta}}$. In other words, the $i$th entry of $\bm{c}_{\bm{\eta}}$ is the variance of the $i$th entry of $\bm{\bar{\sigma}}$ given PMEM parameters $\bm{\eta}$.
@@
    Calculate the posterior mean vector, $\bm{\mu}_{D^{\prime}}$, and the posterior precision vector, $\bm{\beta}_{D^{\prime}}$, for each $D^{\prime}\in\{1,\ldots,D\}$ using Eqs. (23) and (24).
@@
    Update the prior mean vector, $\bm{\eta}$, and the prior precision vector, $\bm{\alpha}$, using Eqs. (28) and (29).
@@ 3.5 Clustering and Markovian approaches
The ELA, or PMEM, is closer to the conventional clustering approach than to the HMM in this aspect. In estimating the PMEM from data, only the equilibrium distribution of the data is fitted, similar to conventional clustering. With the PMEM, one does not fit the state transition rate from each particular state to another. The state transition rates in the case of ELA are determined by the detailed balance condition (i.e., Eq. (18)) and the structure of the estimated disconnectivity graph.
@@ 4 Software
The energy landscape analysis toolbox (ELAT), accompanying a review paper [54], is a MATLAB package with a graphical user interface [87]. It can estimate the PMEM from data, compute the accuracy of fit, basin graphs in two and three dimensions (the three-dimensional version shows the energy value of each activity pattern on the vertical axis), disconnectivity graph, list of activity patterns at each local minimum (each being an $N$-dimensional binary vector, as in the gray-white vectors shown in Fig. 7), time series of the activity pattern, time series of the basin ID to which the activity pattern at each time belongs, and some features of the constructed energy landscape. The code accompanying an MEG study [39] is provided in MATLAB [88]. It uses ELAT and provides analyses complementary to ELAT such as calculation of the best threshold for binarization and a boostraping for statistical analysis of energy values. A Python implementation of ELAT is available [89] as an accompaniment of [43]. Code for carrying out ELA is also available by other authors in R [90] and Mathematica [91], as an accompaniment of [44]. The code accompanying an EEG study [36] is also written R and capable of estimating the PMEM by pseudo-likelihood maximization and carrying out ELA [92].

### v2

@@ 2.7 Features of energy landscapes
The disconnectivity graph and basin graph are visualizations of the energy landscape. They are useful for us to intuitively understand the dynamics generating the multivariate time series data and differences between different data or between different groups of data (e.g., between a group of healthy individuals and a group of those with disease) in terms of the energy landscape. However, we need to quantify such differences to draw conclusions. We refer to such a quantity as a feature of the energy landscape. A primary usage of a feature is to look for association between the feature and a property of interest in the research domain. For example, a schizophrenia researcher may be interested in seeking correlation between the severity of schizophrenia and a feature of the energy landscape. As another example, one may want to measure the effect of treatment in terms of the change in the value of a feature, especially if the feature is known to correlate with the severity of the disease.
@@
    A variant of this feature is an average over different local minima. In [32], the authors used the average of branch length over all the major local minima in the pruned disconnectivity graph. By the pruned disconnectivity graph, we mean one created by sequentially removing insignificant local minima, to keep only the major local minima (see Section 3.2.2 for the procedure). The branch length for local minimum $\bm{\alpha}_{i}$ is $\min_{j\text{ such that }j\neq i}\{\overline{E}_{i,j}-E(\bm{\alpha}_{i})\}$, where $\overline{E}_{i,j}$ now refers to the smallest energy at which the $i$th and $j$th branches of the pruned disconnectivity graph merge. For example, in Fig 3(e), suppose that the pruning process discards $\bm{\alpha}_{2}$, $\bm{\alpha}_{3}$, and $\bm{\alpha}_{6}$, leaving $\bm{\alpha}_{1}$, $\bm{\alpha}_{4}$, and $\bm{\alpha}_{5}$ as major local minima. This pruning is equivalent to setting the threshold energy to the one shown by the dotted horizontal line in the figure to define three major states, each of which is represented by a major local minimum. Then, the branch length for $\bm{\alpha}_{1}$, $\bm{\alpha}_{4}$, and $\bm{\alpha}_{5}$ are equal to $\overline{E}_{1,4}-E(\bm{\alpha}_{1})$, $\overline{E}_{1,4}-E(\bm{\alpha}_{4})$, and $\overline{E}_{1,5}-E(\bm{\alpha}_{5})$, respectively. The last one is equal to $\overline{E}_{4,5}-E(\bm{\alpha}_{5})$ because $\bm{\alpha}_{1}$ and $\bm{\alpha}_{4}$ merge earlier than $\bm{\alpha}_{5}$ does as we climb up the disconnectivity graph from a low energy value. The average branch length used in [32] is the average of these three branch lengths and represents how distinct the different major local minima are from each other.
@@ 3.1 Phase diagram methods
Fig 10: Schematic of the phase diagram of the SK model. A more common parameter in physics literature is the temperature. However, in this figure, we chose the average and standard deviation of $J_{ij}$ because their interpretation is easier in our context.
The statistical mechanical theory exploited here is that of prototypical spin systems, such as the Sherrington-Kirkpatrick (SK) model [67]. The SK model is a type of Ising model and defined by drawing each $J_{ij}$ $(=J_{ji};1\leq i\neq j\leq N)$ independently from the Gaussian distribution with a given mean and standard deviation. The SK model, as well as other spin system models, shows different phases depending on the parameter values (see Fig 10 for a schematic). In the paramagnetic phase, the spins (i.e., $\sigma_{i}$s) tend to be random and orderless in the sense that the mean activity
@@
Criticality, or phase transition, i.e., the boundary between (typically) the paramagnetic phase and the spin-glass phase in a parameter space (see Fig 10), is of special interest due to rich dynamical behavior of the spin system near there. On these grounds, Ezaki et al. developed a data-driven method to locate a given multivariate time series data as a point in the phase diagram to tell which phase the given data resides in, and quantify how close the data is from the boundary with a different phase [19].
@@
where $\hat{\mu}$, for example, is the value of $\mu$ for the originally estimated PMEM. With Eq (17), if we arbitrarily choose the values of $\mu$ and $s$, then we obtain $\bm{J}$, which specifies one PMEM. For each $\bm{J}$ specified by a $(\mu,s)$ pair, we perform Monte Carlo simulations to produce $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ and calculate observables such as $m$ and $q$. In this manner, we can generate a phase diagram for each observable in terms of $\mu$ and $s$, as in Fig 10.
@@ 3.2.2 Randomization
In Fig 3, local minimum $\bm{\alpha}_{2}$ is shallow in the sense that the energy barrier between $\bm{\alpha}_{2}$ and $\bm{\alpha}_{3}$ is small. Therefore, the dwell time in $\bm{\alpha}_{2}$ would be small. Therefore, although $\bm{\alpha}_{2}$ is technically a local minimum, it may be better to ignore it in practice. Only the local minima that have sufficiently high energy barriers with others are trustable. We already discussed this topic; by choosing a threshold energy level, shown by the dotted line in Fig 3(e), one can classify the local minima into groups. However, how should we choose the threshold? How can one tell whether or not a local minimum is too shallow to be seriously considered? This decision has been rather arbitrary in most studies using the ELA. The root of the problem is basically the same as an often asked problem for clustering algorithms for data: how to determine the number of clusters?
@@
We compute the maximal branch length for a sufficiently large number, $c$, of completely random multivariate time series data. We set $c=10^{2}$ in [32] and use $c=10^{3}$ in the following running example in this section. We denote the mean and standard deviation of the maximal branch length over the $c$ random disconnectivity graphs by $\langle L\rangle$ and $\text{std}(L)$, respectively. Then, we remove the local minimum with the shortest branch length in the original disconnectivity graph as being insignificant if its branch is shorter than $\langle L\rangle+2\times\text{std}(L)$. For example, Fig 7(b) shows the disconnectivity graph for a random multivariate binary time series having the same $N$ ($=7$) and $t_{\max}$ ($=3,272$) as those for the original data whose disconnectivity graph is shown in Fig 7(a). Due to the randomization, the disconnectivity graph in Fig 7(b) has shorter branch lengths than the disconnectivity graph in Fig 7(a) in general. The maximal branch length in Fig 7(b) is that of $\bm{\alpha}^{\prime}_{3}$ and equal to $L=0.125$. Based on $c=10^{3}$ random disconnectivity graphs, we obtained $\langle L\rangle=0.0902$ and $\text{std}(L)=0.0471$. Therefore, $\langle L\rangle+2\times\text{std}(L)=0.184$. The shortest branch of the empirical disconnectivity graph shown in Fig 7(a), which is for $\bm{\alpha}_{3}$, is shorter than $0.184$. Therefore, we regard $\bm{\alpha}_{3}$ as an insignificant local minimum and prune it away.
@@
If we have removed a local minimum according to this procedure, then we recompute the branch length of the local minima affected by the removal. In Fig 7(a), removal of the branch of $\bm{\alpha}_{3}$ does not affect the branch length of any other local minima, especially because $\bm{\alpha}_{4}$ and $\bm{\alpha}_{5}$ merge before they merge $\bm{\alpha}_{3}$. Therefore, there is no need to recompute any branch length. After the possible recomputation, we repeat the above procedure. In other words, if the shortest branch is shorter than $\langle L\rangle+2\times\text{std}(L)$, we remove the branch and recompute the affected branch lengths. For example, in Fig 7(a), we next remove the branch of $\bm{\alpha}_{2}$ because it is shorter than the threshold (i.e., $0.184$). This removal changes the branch length for $\bm{\alpha}_{1}$ from 0.814 to 1.082. Next, we remove the branch of $\bm{\alpha}_{4}$, which increases the branch length for $\bm{\alpha}_{5}$ from $0.761$ to 0.902.
@@
We repeat the procedure until the branches of all the surviving local minima are longer than the threshold in the pruned disconnectivity graph. We refer to the surviving local minima as major local minima. For example, for Fig 7(a), the final, pruned discoonnectivity graph, shown in Fig 7(c), only consists of $\bm{\alpha}_{1}$ and $\bm{\alpha}_{5}$, which are the two major local minima. This pruning algorithm also tells us how to associate each removed local minimum and its basin to a major state. In Fig 7(a), $\bm{\alpha}_{1}$, $\bm{\alpha}_{2}$, and their basins constitute a major state represented by $\bm{\alpha}_{1}$. Similarly, $\bm{\alpha}_{3}$, $\bm{\alpha}_{4}$, $\bm{\alpha}_{5}$, and their basins constitute a major state represented by $\bm{\alpha}_{5}$.
@@ 3.2.3 Random-walk simulations
where $T_{\bm{\sigma}\to\bm{\sigma}^{\prime}}$ is the transition probability of the random walk from $\bm{\sigma}$ to $\bm{\sigma}^{\prime}$. This detailed balance condition is a quantitative formulation of the already made claim that the movement of the state uphill on the energy landscape is not much likely to occur, especially when the energy barrier to be overcome to go to a different activity pattern is high. In this way, moves from a high energy activity pattern to a low energy one are more likely than vice versa. However, the random walk is a stochastic process, so moves from a lower energy activity pattern to a higher one do occur with some probability. Note that Eq (18) is consistent with the stationary density given by Eq (6).
@@ 3.3 Challenges
First, ELA is data-hungry. If we aim at maintaining a high accuracy of fit of the PMEM, we roughly need to increase the data length, $t_{\max}$, twice as we add one variable (i.e., as we increase $N$ by $1$). Therefore, $t_{\max}$ needed increases exponentially in terms of $N$, i.e., in proportion to $2^{N}$ [54]. Then, if we seriously take the accuracy, one cannot go beyond approximately $N=10$ or $N=15$, depending on $t_{\max}$ of the data, while ignoring the fitting accuracy to conduct a large-scale ELA (e.g., see Section 3.1) is a methodological choice.
@@ 3.4 Variational Bayes approximation method
is an $M\times M$ matrix, and $\rm{diag}(\cdot)$ represents the diagonal matrix whose entries are given by the arguments. In Eq (23),
@@
$\bm{\bar{\sigma}}\equiv(\sigma_{1},\ldots,\sigma_{N},\sigma_{1}\sigma_{2},\sigma_{1}\sigma_{3},\ldots,\sigma_{N-1}\sigma_{N})^{\top}$ (27)
when the PMEM parameters $(h_{1},\ldots,h_{N},J_{12},J_{13},\ldots,J_{N-1,N})$ are given by $\bm{\eta}$; matrix $\rm C_{\bm{\eta}}$ is the $M\times M$ covariance matrix of $\bm{\bar{\sigma}}$ when the PMEM parameters are $\bm{\eta}$. In Eq (24), $\bm{c}_{\bm{\eta}}$ is the vector composed of the diagonal entries of $\rm C_{\bm{\eta}}$. In other words, the $i$th entry of $\bm{c}_{\bm{\eta}}$ is the variance of the $i$th entry of $\bm{\bar{\sigma}}$ given PMEM parameters $\bm{\eta}$.
@@
    Calculate the posterior mean vector, $\bm{\mu}_{D^{\prime}}$, and the posterior precision vector, $\bm{\beta}_{D^{\prime}}$, for each $D^{\prime}\in\{1,\ldots,D\}$ using Eqs (23) and (24).
@@
    Update the prior mean vector, $\bm{\eta}$, and the prior precision vector, $\bm{\alpha}$, using Eqs (28) and (29).
@@ 3.5 Clustering and Markovian approaches
The ELA, or PMEM, is closer to the conventional clustering approach than to the HMM in this aspect. In estimating the PMEM from data, only the equilibrium distribution of the data is fitted, similar to conventional clustering. With the PMEM, one does not fit the state transition rate from each particular state to another. The state transition rates in the case of ELA are determined by the detailed balance condition (i.e., Eq (18)) and the structure of the estimated disconnectivity graph.
@@ 4 Software
The energy landscape analysis toolbox (ELAT), accompanying a review paper [54], is a MATLAB package with a graphical user interface [87]. It can estimate the PMEM from data, compute the accuracy of fit, basin graphs in two and three dimensions (the three-dimensional version shows the energy value of each activity pattern on the vertical axis), disconnectivity graph, list of activity patterns at each local minimum (each being an $N$-dimensional binary vector, as in the gray-white vectors shown in Fig 7), time series of the activity pattern, time series of the basin ID to which the activity pattern at each time belongs, and some features of the constructed energy landscape. The code accompanying an MEG study [39] is provided in MATLAB [88]. It uses ELAT and provides analyses complementary to ELAT such as calculation of the best threshold for binarization and a boostraping for statistical analysis of energy values. A Python implementation of ELAT is available [89] as an accompaniment of [43]. Code for carrying out ELA is also available by other authors in R [90] and Mathematica [91], as an accompaniment of [44]. The code accompanying an EEG study [36] is also written R and capable of estimating the PMEM by pseudo-likelihood maximization and carrying out ELA [92].